If the average time spent studying is 10 hours and the standard deviation is 4 hours, what it the probability that a student will spend more than 13 hours studying?


1
Expert's answer
2022-06-02T18:19:35-0400
"P(X>13)=1-P(Z\le\dfrac{13-10}{4})"

"=1-P(Z\le0.75)\approx0.2266"